OpenStudy (anonymous):

A triangle has side lengths of 15,20, and 25. What type of triangle is it?

OpenStudy (anonymous):

scalene as all sides r of different lengths

OpenStudy (anonymous):

it will also be a right angled triangle because (15)^2 + (20)^2 = 25^2 225 + 400 = 625

OpenStudy (anonymous):

acute,obtuse, right ,equiangular?

OpenStudy (anonymous):

as i just wrote above, it will b a right triangle